OpenStudy (thatonegirl_):

Find an equation of the line that is tangent to the graph of f and parallel to the given line..

OpenStudy (thatonegirl_):

\[f(x)=3x^3\] \[y=9x+9\]

OpenStudy (campbell_st):

ok... so I'd expect this is from a calculus course. you need to find the point of contact between the tangent and curve you know the slope of the tangent is 9. so differentiate f(x) . This will give the equation for the slope of the tangent at any point on the curve. then let f'(x) = 9 and solve for x. substitute this value into the original function to find f(x) then you will have a point (x, f(x)) and a slope m = 9 you will then be able to find the equation of the tangent. Hope it makes sense.

OpenStudy (ajspeller):

\[f'(x)=9x^2\] Solve\[9=9x^2\] use the point (x,f(x)) with slope, m=9 into the point slope formula \[(y-f(x))=m(x-x) \] where the second x is the x from (x,f(x))

OpenStudy (thatonegirl_):

Thank you both!